# Quenched shear stresses and shear moduli in disordered elastic networks

### v1

@@ Introduction.
Equilibrium CM-distributions.
Let us first characterize the moments for switched-off reorganization MC moves ($\omega=0$). Using a half-logarithmic representation this is shown in Fig. 2 with the sampling time $t$ being the horizontal axis. As expected, all moments become rapidly $t$-independent ($t\gg 1$) as emphasized by horizontal lines. We note especially the static asymptotic limits
@@
The inequalities Eq. (8) and Eq. (10) thus hold as expected. We also emphasize that $\mu_{\mathrm{a}}$ is very small albeit finite at variance to liquid systems, cf. Eq. (11), and the $c$-averaged shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}\approx\mu_{1}$ is thus dominated by the variance $\mu_{1}$ of the quenched stresses.
@@
The thermodynamic static values for permanent CMs are also of relevance if the reorganization MC moves are not completely switched off. This is demonstrated for a large range of $\omega$ in Fig. 3 for $\mu(t)$. As can be seen, $\mu(t)\to\mu_{\mathrm{A}}$ for all $\omega$ for very small times (thin horizontal line) foo (c). This is expected since $\mu_{\mathrm{F}}(t)$ must vanish in this limit. For small $\omega$ and not too large $t$, $\mu(t)$ approaches the static shear modulus $\mu_{\mathrm{q}}$ of quenched networks (bold horizontal line). $\mu_{\mathrm{q}}$ becomes an intermediate upper shoulder (plateau) upon further increasing $\omega$ or $t$. Importantly, for sufficiently large $t$ and $\omega$ the dynamical shear modulus $\mu(t)$ always approaches $\mu_{\mathrm{a}}$ (dashed horizontal line), i.e. the static shear modulus of completely annealed networks. Naturally, this is best seen for our largest reorganization frequencies ($\omega\gg 0.01$). We note finally that the relaxation time $t_{\star}$, operationally defined by $\mu(t=t_{\star})/\mu_{\mathrm{a}}-1=5\%$, increases inversely with $\omega$.
Out-of-equilibrium CM-distributions.
Having confirmed Eq. (10) for equilibrium CM-distributions we show now using two examples with non-equilibrium distributions that these inequalities do not hold in general.
@@
Figure 4: Time-dependence of observables after quenching at $t=0$ from $T=10$ to $T=0.1$ and allowing the CMs to reorganize with a small frequency $\omega=0.01$. Eq. (10) gets violated for $t\ll t_{\star}\approx 30000$ (vertical line) where $\mu_{1}(t)$, $\mu_{2}(t)$ and $-\mu_{\mathrm{a}}(t)$ are found to decay logarithmically.
@@
Starting with annealed liquids at high temperatures amorphous solids are commonly created by a rapid quench below the glass transition temperature $T_{\mathrm{g}}$ Götze (2009). This has not only the effect that the relaxation dynamics is strongly reduced but, moreover, that glasses are generally not at thermal equilibrium at $T<T_{\mathrm{g}}$, i.e. Eq. (10) may not hold. We consider here a similar situation where our networks are first equilibrated with finite $\omega$ at $T=10$ and then suddenly at $t=0$ quenched to our standard temperature $T=0.1$, i.e. the CM-distribution is initially not at equilibrium. As shown in Fig. 4, it is then allowed to relax with $\omega=0.01$. The data points have been obtained using an iteration of subsequent tempering and measuring intervals of both $\Delta t=1000$. The indicated time for the horizontal axis is $t=\Delta t(2i-0.5)$ for each iteration step $i$. Interestingly, Eq. (10) is strongly violated for $t\ll t_{\star}\approx 30000$. For larger $t$ the equilibrium CM-distribution for $T=0.1$ is restored, cf. Eq. (12).
@@
Figure 5: Scalar active two-temperature model with $T=0.1$ for the network vertices and $T_{\mathrm{CM}}$ for the CMs. A half-logarithmic representation of the observables as a function of the reduced temperature $x=T_{\mathrm{CM}}/T$ is used. The final production runs are performed at switched-off MC moves ($\omega=0$) sampling over a sampling time $t=10^{4}$. Eq. (10) is violated for $x\gg 1$.
@@
As a second example we present in Fig. 5 data for two-temperature networks where the MD dynamics of the vertices is coupled to a Langevin thermostat with $T=0.1$ while a broad range of temperatures $T_{\mathrm{CM}}$ is imposed for the MC steps of the network reorganization. This model is motivated by recent work on scalar active matter with different temperatures for each particle class Grosberg and Joanny (2015); Elismaili et al. (2020). For each $x=T_{\mathrm{CM}}/T$ we anneal the CM-ensemble using a finite $\omega$ over sufficiently long times until all properties become stationary. The reorganization MC steps are then switched off for the production runs presented in Fig. 5. As expected, the data points for $x\approx 1$ are consistent with Fig. 2 and $\mu_{\mathrm{A}}(x)$, $\mu_{\mathrm{F}}(x)$ and $\mu_{\mathrm{q}}(x)$ are seen to depend very weakly on $x$. Importantly, we observe for $x\gg 1$ that the inequalities Eq. (10) are strikingly violated, just as for the short-time limit in Fig. 4. This makes sense since in both cases the CMs are distributed using a too large temperature which generates stronger quenched stresses. Moreover, Eq. (10) does not only hold for $x\approx 1$ but also for $x\ll 1$. We note finally that in the small-$x$ limit the moments $\mu_{1}(x)$ and $\mu_{2}(x)$, measuring the quenched shear stresses, slightly decrease albeit remaining finite. Hence, $-\mu_{\mathrm{a}}(x)=\mu_{2}(x)-\mu_{\mathrm{A}}(x)$ also decreases with $-\mu_{\mathrm{a}}(x)\to-3$ for $x\to 0^{+}$.
Conclusion.
We have considered in this work various contributions to the shear modulus of disordered elastic networks with annealed or quenched CMs. By rewriting the relevant relation of the stress-fluctuation formalism it was shown that the quenched shear stresses induced by the spontaneous symmetry breaking with respect to Eq. (1) yield an important contribution $\mu_{1}$ to the shear modulus $\mu_{\mathrm{q}}=\mu_{1}+\mu_{\mathrm{a}}$. The second contribution $\mu_{\mathrm{a}}$ is a simple average over all states and CMs, cf. Eq. (9), which for equilibrium CM-distributions is identical to the shear modulus of annealed networks (cf. Fig. 3). Interestingly, $\mu_{\mathrm{a}}$ may be strongly negative for out-of-equilibrium distributions as shown by considering a temperature quench (cf. Fig. 4) and a scalar active two-temperature model (cf. Fig. 5). While the shear modulus $\mu_{\mathrm{q}}$ can still be computed in these cases using the stress-fluctuation relation Eq. (7) the more special inequalities Eq. (10) do not hold.
The presented relations should also be relevant for amorphous colloidal or polymer glasses formed by a temperature quench from a liquid phase. For an extremely slow quench protocol such system may stay at equilibrium and according to Eq. (11) the static shear modulus $\mu_{\mathrm{q}}=\mu_{1}$ is set by the variance of the quenched shear stresses. More generally, however, the interacting particles may not be at equilibrium and Eq. (10) and Eq. (11) may not hold. Preliminary numerical work for colloidal glasses obtained using either rapid $T$-jumps or more gradual $T$-quenches shows indeed that $\mu_{\mathrm{a}}(T)$ becomes strongly negative below $T_{\mathrm{g}}$.
@@ Acknowledgments.
We are indebted to J. Farago and J. Baschnagel (Strasbourg) for helpful discussions.

### v2

@@ IV Elastic networks at equilibrium
Quenched CM-distributions.
Let us first characterize the moments for switched-off reorganization MC moves ($\omega=0$, $t_{\star}=\infty$). Using a half-logarithmic representation this is shown in Fig. 2 with the sampling time $t$ being the horizontal axis. As expected, all moments become rapidly $t$-independent ($t\gg 1$) as emphasized by horizontal lines. The static asymptotic limits are
@@
The inequalities Eq. (12) and Eq. (14) thus hold as expected. Please note that $\mu_{\mathrm{a}}$ is very small albeit finite.
@@
The thermodynamic static values for permanent CMs are also of relevance if the reorganization MC moves are not completely switched off. This is demonstrated for a large range of $\omega$ in Fig. 3 for $\mu(t)$. As can be seen, $\mu(t)\to\mu_{\mathrm{A}}$ for all $\omega>0$ for very small times (thin horizontal line) foo (c). This is expected since $\mu_{\mathrm{F}}(t)$ must vanish in this limit. For small $\omega$ and not too large $t$, $\mu(t)$ approaches the static shear modulus $\mu_{\mathrm{q}}$ of quenched networks (bold horizontal line). $\mu_{\mathrm{q}}$ becomes an intermediate upper shoulder (plateau) upon further increasing $\omega$ or $t$. Importantly, for sufficiently large $t$ and $\omega$ the dynamical shear modulus $\mu(t)$ always approaches $\mu_{\mathrm{a}}$ (dashed horizontal line), i.e. the static shear modulus of completely annealed networks. Naturally, this is best seen for our largest reorganization frequencies ($\omega\gg 0.01$). We note finally that the relaxation time $t_{\star}$, operationally defined by $\mu(t=t_{\star})/\mu_{\mathrm{a}}-1=5\%$, is inversely proportional to $\omega$ and that it is possible (not shown) to collapse the data for large $t$ by tracing $\mu(t)$ as a function of $t/t_{\star}(\omega)$.
V Out-of-equilibrium networks
@@
Figure 4: Time-dependence of various moments after quenching at $t=0$ from $T=10$ to $T=0.1$ and allowing the CMs to reorganize with a small frequency $\omega=0.01$. Eq. (14) gets violated for $t\ll t_{\star}\approx 30000$ (vertical line) where $\mu_{1}(t)$, $\mu_{2}(t)$ and $-\mu_{\mathrm{a}}(t)$ are found to decay logarithmically.
Having confirmed Eq. (14) for equilibrium CMs we show now using two examples with non-equilibrium CMs that these inequalities do not hold in general.
Temperature jump.
@@
Starting with annealed liquids at high temperatures amorphous solids are commonly created by a rapid quench below the glass transition temperature $T_{\mathrm{g}}$ Dhont (2001). This has not only the effect that the relaxation dynamics is strongly reduced but, moreover, that glasses are generally not at thermal equilibrium at $T<T_{\mathrm{g}}$, i.e. Eq. (14) may not hold. This issue will be addressed in Sec. VI. Here we consider first a similar situation where networks are first equilibrated with finite $\omega$ at $T=10$ and then suddenly at $t=0$ quenched to our standard temperature $T=0.1$, i.e. the CM-distribution is initially not at equilibrium. As shown in Fig. 4, it is then allowed to relax with $\omega=0.01$. The data points have been obtained using an iteration of subsequent tempering and measuring intervals of both $\Delta t=1000$. The indicated time for the horizontal axis is $t=\Delta t(2i-0.5)$ for each iteration step $i$. While Eq. (12) always holds it is seen that Eq. (14) is strongly violated for short times $t\ll t_{\star}$ with $t_{\star}\approx 30000$. In the large-$t$ limit the equilibrium CM-distribution for $T=0.1$ is restored, cf. Eq. (15).
@@
Figure 5: Scalar active two-temperature model with $T=0.1$ for the network vertices and $T_{\mathrm{CM}}$ for the CMs. A half-logarithmic representation of the observables as a function of the reduced temperature $x=T_{\mathrm{CM}}/T$ is used. The final production runs are performed at switched-off MC moves ($\omega=0$) sampling over a sampling time $t=10^{4}$. Equation (14) is violated for $x\gg 1$ but holds for $x\ll 1$.
Active two-temperature model.
@@
As a second example we present in Fig. 5 data for two-temperature networks where the MD dynamics of the vertices is coupled to a Langevin thermostat with $T=0.1$ while a broad range of temperatures $T_{\mathrm{CM}}$ is imposed for the MC steps of the network reorganization. This model is motivated by recent work on scalar active matter with different temperatures for each particle class Grosberg and Joanny (2015); Elismaili et al. (2020). Half-logarithmic coordinates are used with $x=T_{\mathrm{CM}}/T$ being the dimensionless horizontal axis. For each $x$ we anneal the CM-ensemble using a finite $\omega$ over sufficiently long times until all properties become stationary. The reorganization MC steps are then switched off for the production runs presented in Fig. 5 ($\omega=0$, $t_{\star}=\infty$, $t=10^{4}$). As expected, the data points for $x\approx 1$ are consistent with Fig. 2 and $\mu_{\mathrm{A}}(x)$, $\mu_{\mathrm{F}}(x)$ and $\mu_{\mathrm{q}}(x)$ are seen to only depend very weakly on $x$. Importantly, we observe for $x\gg 1$ that the inequalities Eq. (14) are strikingly violated, just as for the short-time limit in Fig. 4. This makes sense since in both cases the CMs are distributed using a too large temperature which generates stronger quenched stresses. Moreover, Eq. (14) does not only hold for $x\approx 1$ but also for $x\ll 1$. We also note that in the small-$x$ limit the moments $\mu_{1}(x)$ and $\mu_{2}(x)$, measuring the quenched shear stresses, slightly decrease albeit remaining finite. Hence, $-\mu_{\mathrm{a}}(x)=\mu_{2}(x)-\mu_{\mathrm{A}}(x)$ also decreases with $-\mu_{\mathrm{a}}(x)\to-3$ for $x\to 0^{+}$.
VI Liquid-type elastic solids
General considerations.
We have considered above elastic networks which by construction must have a finite shear modulus even for annealed reference lengths $R_{s}$ where $\mu=\mu_{\mathrm{a}}>0$. We turn now to liquid-type systems such as reversible physical gels Wittmer et al. (2016); Bouzid and Gado (2018); Zhang et al. (2019) or polymer melts Rubinstein and Colby (2003) where
$\mu\simeq\mu_{\mathrm{a}}\simeq\mu_{1}\simeq 0\mbox{ for }t\gg t_{\star}$ (16)
with $t_{\star}$ being the finite terminal relaxation time of the system. Obviously, in many real experiments and computer simulations $t_{\star}$ becomes rapidly extremely large, especially upon cooling, and in practice $t\ll t_{\star}$ for all realistic sampling times $t$. One thus expects elastic solid behavior with a finite shear modulus $\mu(t)\approx\mu_{\mathrm{q}}>0$ for a finite $t$-window similar to the low-$\omega$ data presented in Fig. 3 for elastic networks. Occasionally, such effectively quenched liquid-type systems may still be at thermal equilibrium Wittmer et al. (2016). (For reversible physical gels this may be achieved, e.g., by temporarily adding some catalytic compounds facilitating the network reorganization or in computational work by means of additional efficient albeit artificial MC moves as shown below for one example.) Assuming thus an equilibrium distribution of dynamically quenched configurations, the inequalities Eq. (14) are expected to become equalities, i.e.
$0=\mu_{\mathrm{a}}\equiv\mu_{\mathrm{A}}-\mu_{2}=\mu_{\mathrm{q}}-\mu_{1}.$ (17)
In agreement with related recent studies Wittmer et al. (2016); Klochko et al. (2021); George et al. (2021) we have thus finally confirmed Eq. (1). Note that failure of Eq. (17) implies that the quenched configuration ensemble is inconsistent with an equilibrium distribution.
Figure 6: $T$-dependence of contributions to the shear modulus for $n=10000$ pLJ particles at an imposed average normal pressure $P=2$ Wittmer et al. (2013); George et al. (2021). All data points are obtained using MD-production runs over $t=10^{5}$ and averaged over $N_{\mathrm{c}}=100$ independent configurations. Open symbols refer to configurations obtained by means of a constant-rate $T$-quench while the stars show $-\mu_{\mathrm{a}}$ for configurations being first thoroughly equilibrated using a mix of local and swap MC moves Ninarello et al. (2017); George et al. (2021). While Eq. (17) is violated in the first case below $T_{\mathrm{g}}\approx 0.26$ (vertical dashed line) it holds in the second case even for $T\ll T_{\mathrm{g}}$ (horizontal dashed line).
Colloidal glass-former.
The above statements are illustrated in Fig. 6 using systems of polydisperse Lennard-Jones (pLJ) particles which have been extensively investigated in the past Wittmer et al. (2002, 2013); Klochko et al. (2021); George et al. (2021). We average over $N_{\mathrm{c}}=100$ independent configurations which are kept at an average normal pressure $P=2$ by means of a standard barostat Allen and Tildesley (2017). All production runs are performed using straight-forward MD (with switched off barostat) just as for the spring networks investigated above. We only present data for $t=10^{5}$. Two different cases of ensemble preparation are compared. The open symbols in Fig. 6 refer to configurations created starting from equilibrated start configurations at $T=0.5$ using a standard quench protocol decreasing the imposed temperature with a constant quench rate $0.5\cdot 10^{-5}$ (in LJ units) and a final tempering over $t=10^{5}$ for each $T$ sampled. In addition, we thoroughly equilibrated each configuration using a mix of local and swap MC moves over $10^{6}$ MC cycles Ninarello et al. (2017); George et al. (2021) and sample this second ensemble again using MD. For clarity, we only show for this second case data for $-\mu_{\mathrm{a}}$ (stars). This shows that for well-equilibrated systems Eq. (17) holds in principle for all temperatures. (As revealed by closer inspection, $\mu_{\mathrm{a}}=0$ is slightly violated for our lowest temperatures suggesting that additional equilibration with swap MC moves is warranted.) This is different for the first data set (open symbols) obtained using the standard MD-quench for which Eq. (17) is clearly violated below the glass-transition temperature $T_{\mathrm{g}}\approx 0.26$ Wittmer et al. (2013); Klochko et al. (2021). Especially, it is seen that $\mu_{1}\gg\mu$ and $-\mu_{\mathrm{a}}=\mu_{2}-\mu_{\mathrm{A}}\gg 0$. We thus found similar behavior as for the networks in Fig. 4.
VII Conclusion
Summary.
We have investigated in this work general inequality relations for the different contributions to the shear modulus of amorphous bodies with annealed or quenched configurations. We compared ensembles of quenched configurations at thermal equilibrium with more general distributions with a different dispersion of quenched shear stresses. To illustrate our key points we have first focused on simple elastic spring networks for which the disordered CMs can be simply defined and operationally changed in terms of the reference lengths $R_{s}$ of the ideal springs, cf. Fig. 1. The stress-fluctuation formalism relation for the shear modulus was rewritten as $\mu=\mu_{1}+\mu_{\mathrm{a}}$ with $\mu_{1}\geq 0$ being the (rescaled) variance of the quenched shear stresses and $\mu_{\mathrm{a}}$ a simple average over all states and CMs, cf. Eq. (13). For annealed networks $\mu_{\mathrm{a}}$ is equivalent to the shear modulus $\mu$ as shown in Fig. 3. Interestingly, $\mu_{\mathrm{a}}$ may be strongly negative for out-of-equilibrium distributions as shown by considering a temperature quench (cf. Fig. 4) and a scalar active two-temperature model (cf. Fig. 5). While the shear modulus can still be computed in these cases using the stress-fluctuation relation Eq. (11), the more special inequalities Eq. (14) do not hold in general and $\mu_{1}$ may thus differ strongly from the shear modulus. Turning in Sec. VI to glass-forming pLJ particles we have confirmed that similar behavior can be expected for other types of amorphous elastic solids. While Eq. (17) was seen to hold for pLJ glasses equilibrated by means of the swap MC algorithm, systematic and strong deviations arise for standard out-of-equilibrium ensembles prepared using a conventional quench protocol (cf. Fig. 6). The measurement of deviations from Eq. (17) for liquid-type systems provides thus quite generally a convenient diagnosis tool of the equilibration status of configuration ensembles.
Outlook.
Further work will focus on the systematic analysis of effects related to the finite sampling time $t$ used for our pLJ systems and the characterization of the equilibration of the configuration ensemble by tempering with different frequencies of swap MC moves. Relaxation behavior similar to Fig. 4 is expected. It may also be of interest to consider scalar active self-assembled networks of colloids bridged by ideal springs Wittmer et al. (2016) and polydisperse particles with different temperatures as in Refs. Grosberg and Joanny (2015); Elismaili et al. (2020). We expect Eq. (17) to be violated just as for the two-temperature active network presented in Fig. 5. The above work has focused on macroscopic observables. We note finally that similar relations should hold for the microscopic shear modulus and its contributions at finite wavevectors $q$ Wittmer et al. (2023); Semenov and Baschnagel (2024). For melts of long reptating polymer chains at equilibrium it is, e.g., expected that Eq. (1) remains valid in reciprocal space for sampling times below the reptation time Rubinstein and Colby (2003). It then follows from general considerations of isotropic tensor fields Wittmer et al. (2023) that the spatial correlations of the (transiently) quenched shear stresses must be long ranged decaying as $1/r^{3}$ with the distance $r$ and the prefactor being set by the intermediate plateau $G(t)\approx\mu$ of the shear stress relaxation function. Work in this direction is currently underway.
@@ Acknowledgments.
We are indebted to J. Baschnagel (Strasbourg) for helpful discussions.